HETAV PATEL

Grade 11,

If the velocity(V),force(F) and energy(E) are taken as fundamental units, then dimensional formula if mass will be _____? help me please

If the velocity(V),force(F) and energy(E) are taken as fundamental units, then dimensional formula if mass will be _____?


help me please

Grade:

2 Answers

vikas askiitian expert
13 years ago

let the dimensional formula of mass is M = EaFbVc

 E=ML2T-2  ,  F =MLT-2  , V=LT-1

on putting these

    M = Ma+b L2a+b+c T-2a-2b-c

equating powers ,we get

 a=1 , b=0, c=-2

 dimensional formula of mass= E1.F0 V-2

Kushagra Madhukar
3 years ago
Dear student,
Please find the solution to your problem.
 
Let the dimensional formula of mass is M = EaFbVc
E = ML2T-2  ,  F = MLT-2  , V = LT-1
on putting these
    M = Ma+b L2a+b+c T-2a-2b-c
equating powers ,we get
a = 1 , b = 0, c = – 2
Hence, dimensional formula of mass = E1.F0 V-2 = E/V2
 
Thanks and regards,
Kushagra